OpenStudy (anonymous):

If f(x) = x + 2, find f(3). 1 6 5 −1

OpenStudy (anonymous):

\[f(3)=3+2\]

OpenStudy (anonymous):

yeah its 5

OpenStudy (anonymous):

just plug it in. f(x), the x value just means the variable you will use each time x pops up in the equation, so you plug it in. in this case, your equation says f(x)=x+2, so you plug 2 in and you have f(3)=3+ 2. So your answer will be 5. just remember, plug it in!